# Region-Wise Correspondence Prediction between Manga Line Art Images

### v2

@@ Introduction
3D-Rendered Line Art. We use the PaintBucket-Character (PBC) (Dai et al. 2024) dataset for both training and evaluation. We sample pairs of frames separated by seven frames within the same animation sequence, defining region correspondences based on shared part-level semantic labels. For example, region 19 in the left image of Figure 5 and region 40 in the right image share label ID 23, indicating a semantic match.
@@
Hand-Drawn Manga Style Line Art. For training on hand-drawn style data, we use an in-house dataset constructed from short anime clips. Frame pairs are sampled every 18 frames, converted to line art using the MangaNinja method (Liu et al. 2025), and annotated using our automatic pipeline, yielding 20,000 ( 39,602 images) image pairs.
For evaluation, we use the ATD-12K (ATD) (Siyao et al. 2021) dataset and a synthetic dataset (GenAI) generated via the Imagen diffusion model (Saharia et al. 2022). For both datasets, we extract line art using the same method and manually refine region correspondences – 25 pairs (50 images) for ATD and 20 pairs (40 images) for GenAI. An example of the annotated data is shown in Figure 4.
Figure 5: Example of region-level correspondence annotations in the PBC dataset. Matched regions are visualized using the same color across two images. Yellow boxes display the region ID and label ID of each region in the format Region ID (Label ID).
Implementation Details
We adopt a ViT-B/16 backbone pre-trained on ImageNet (Dosovitskiy et al. 2020), and normalize the input images using the standard ImageNet mean and standard deviation to match the requirements of the pre-trained model. To fully leverage the capacity of ViT, we use its default patch size of $16\times 16$. Our Multiplex Transformer consists of $M=4$ stacked layers of self-attention and cross-attention. The model is trained for 20 epochs with a batch size of 16, using the AdamW optimizer and an initial learning rate of $1\times 10^{-4}$. A learning rate scheduler combining a warm-up phase with cosine annealing is applied. The whole process is conducted on a single NVIDIA A100 GPU.
Quantitative Results
Datasets Matching Type AP Best F1 Precision Recall Top-1 Accuracy Top-5 Accuracy
PBC Intra-img ($I_{a}$) 99.92 98.57 99.06 98.08
@@
ATD Intra-img ($I_{a}$) 85.40 76.79 78.16 75.46
Intra-img ($I_{b}$) 82.11 73.18 69.70 77.03
Cross-img 80.53 71.71 69.10 74.53 83.59 92.17
GenAI Intra-img ($I_{a}$) 84.52 76.28 74.89 77.73
Intra-img ($I_{b}$) 86.22 77.34 74.36 80.57
Cross-img 84.48 75.52 72.23 79.14 75.94 85.74
Table 1: Patch-level evaluation on PBC, ATD, and GenAI datasets (percentage values).
Dataset Matching Type ARI mIoU (P→G) mIoU (G→P) CR Region Precision Region Recall
PBC Intra-img ($I_{a}$) 83.05 60.20 20.30 0.32
@@
Cross-img 59.92 21.37
ATD Intra-img ($I_{a}$) 55.11 32.37 37.88 1.37
Intra-img ($I_{b}$) 50.37 29.74 35.90 1.49
Cross-img 72.41 32.93
GenAI Intra-img ($I_{a}$) 46.82 37.07 31.59 1.04
Intra-img ($I_{b}$) 56.81 40.35 36.38 1.07
Cross-img 68.52 22.05
@@
Table 2: Region-level evaluation on PBC, ATD, and GenAI datasets (percentage values; CR is shown in raw value). Region Precision and Region Recall are computed using predicted region pairs with purity >> 0.8.
Patch-Level Evaluation. We first assess how well our Transformer-based model captures patch-level semantic similarities by comparing the predicted similarity matrix $S$ against the ground-truth correspondence matrix $G$. We conduct experiments on three datasets: PBC, ATD, and GenAI, and report the results in Table 1.
We evaluate both intra-image similarity (within the same image) and cross-image similarity (between image pairs) using the following metrics. First, we compute precision-recall (PR) curves (Figure 6) by thresholding similarity scores and report standard metrics including Average Precision (AP), best F1 score, and the corresponding precision and recall at the optimal threshold. For cross-image evaluation, we also report top-$K$ accuracy, defined as the percentage of ground-truth matched patches whose true match appears among the top-$K$ most similar patches in the other image.
As shown in Table 1 and Figure 6, our model achieves near-perfect intra-image performance on the PBC dataset, with AP scores above 99.9% and best F1 scores over 98.5% for both $I_{a}$ and $I_{b}$. This indicates that the model effectively captures internal region consistency in well-structured synthetic line art. On the more challenging ATD and GenAI datasets-designed to simulate real-world line art-intra-image AP remains above 82%, and best F1 scores exceed 73%, demonstrating the model’s strong generalization.
In the cross-image setting, our model achieves a level of performance comparable to the intra-image case, confirming its ability to learn patch-level correspondences across image pairs. This robustness is largely attributed to our Transformer’s joint learning of intra-image and cross-image features, which enables effective matching even in the absence of color and texture cues.
Figure 6: Precision-recall (PR) curves for intra-image and cross-image patch-level matching.
Region-Level Evaluation. We further evaluate our method at the region level, focusing on two aspects: (1) intra-image region grouping via patch merging, and (2) cross-image region correspondence prediction. Results on the PBC, ATD, and GenAI datasets are shown in Table 2.
@@
Figure 7: Visualization examples on the PBC and GenAI datasets. In the region segmentation results, the same color within an image denotes a single region, and white indicates the background. In the region matching results, the same color across two images indicates a matched region pair, while gray denotes regions without a match.
To assess the quality of patch clustering within each image, we compare the predicted region groups $\mathcal{R}_{a}$ and $\mathcal{R}_{b}$ with the pixel-level ground-truth segmentation using three metrics: Adjusted Rand Index (ARI), mean Intersection over Union from prediction to ground truth (mIoU P→G), and from ground truth to prediction (mIoU G→P). We also report the Cluster Ratio (CR), which measures the average number of predicted regions per ground-truth region. Ideally, CR should be close to 1, indicating a balanced segmentation. A CR much lower than 1 suggests under-segmentation (i.e., multiple ground-truth regions are merged into one), while a CR significantly higher than 1 implies over-segmentation, where individual regions are unnecessarily fragmented. For cross-image evaluation, we adopt a strategy inspired by visual relationship detection (Xu et al. 2017). Specifically, we compute region-wise precision and recall based on region pairs whose predicted correspondences have a purity score greater than 0.8. This ensures that only meaningful and consistent matches are considered.
As shown in Table 2, our post-processing pipeline performs well on the PBC dataset, which features clean, closed contours and fewer regions per image. Specifically, we achieve an ARI exceeding 83% and an mIoU (P→G) of approximately 60%, indicating robust intra-image clustering performance. In contrast, performance on ATD and GenAI is lower, reflecting the challenges posed by real-world or realistically styled line art,. In terms of cross-image region matching, our method achieves reasonable region precision, demonstrating its ability to identify meaningful correspondences. However, region recall remains relatively low (20-30%), suggesting that our method struggles with matching small fragmented regions or large semantic regions that are inconsistently segmented (e.g., hair or clothing divided into multiple parts). This highlights the difficulty of consistent region matching in highly abstract line art.
Qualitative Results
@@
We qualitatively evaluate our method on multiple datasets, with examples shown in Figure 7. As in our quantitative evaluation, we analyze the results from two aspects: (1) intra-image region segmentation and (2) cross-image region matching.
@@
For region segmentation, we compare our method with the Segment Anything Model (SAM) (Kirillov et al. 2023), a strong baseline for natural images. As shown in Figure 7, SAM often merges distinct semantic parts (e.g., hair or clothing) with the background and fails to produce fine-grained boundaries due to the lack of color, texture, and shading cues. In contrast, our method generates more precise and semantically coherent regions.
For region matching, our approach accurately aligns major regions, including challenging PBC scenes with large pose changes and significant shape deformation. Our region correspondence results demonstrate promising performance and align well with human perception in many cases. This is inherently challenging and has received limited prior attention, with no established baselines for direct comparison. To our knowledge, our method is the first to enable such correspondence for manga line art.
Limitation
Nevertheless, there is still a gap compared with manually annotated ground truth for cross-image region matching. For instance, to avoid over-segmentation, we merge small regions during post-processing. This can lead to the loss of fine-grained semantic details – such as eyes and eyebrows being merged into larger facial regions. We acknowledge this trade-off as a current limitation and plan to address it through improvements to both the model architecture and the post-processing pipeline. Additionally, we aim to extend our approach to handle more diverse manga styles and complex scenes, improve its robustness in real-world scenarios, and explore its integration into downstream applications.
Conclusion
In this paper, we explored a realistic and previously underexplored task: predicting region-wise correspondence between manga line art images without any prior segmentation or annotations. We proposed a Transformer-based framework that jointly learns intra-image and cross-image features from raw line art images, and introduced a post-processing method that aggregates and refines the patch-level predictions into region-level correspondences. To support training and evaluation, we developed an automatic annotation pipeline and manually refined a subset of the annotations to construct benchmark evaluation datasets. Experiments on multiple datasets demonstrate that our model achieves high patch-level accuracy and produces coherent region-level correspondences. In future work, we plan to apply our region correspondence model to downstream tasks such as automatic line art colorization and frame interpolation, where consistent region-wise understanding plays a critical role.

### v3

@@ B.1 Dataset
The dataset contains over 10,000 synthetic line art images paired with part-level semantic labels under consistent character poses and viewpoints. We sample line art pairs by selecting frames separated by seven frames within the same animation sequence, and define region correspondences based on shared semantic part labels. For example, region 19 in the left image of Figure A and region 40 in the right image share label ID 23, indicating a semantic correspondence.
@@
B.2 Quantitative Results
Patch-level and region-level evaluation results on the PBC dataset are reported in Table C and Table D, respectively.
As shown in Table C, our model achieves near-perfect intra-image patch matching performance, with AP scores above 99.9% and best F1 scores exceeding 98.5% for both $I_{a}$ and $I_{b}$. This demonstrates that the model effectively captures internal region consistency in well-structured synthetic line art.
At the region level, we tune post-processing parameters to preserve visually coherent segmentation while maximizing prediction accuracy. As shown in Table D, our method performs well on both region segmentation and region correspondence, achieving a correspondence accuracy of 59.9%. However, because the PBC dataset contains many small and fine-grained regions, the Cluster Ratio (CR) and region recall remain relatively low, typically around 30%. While reducing over-segmentation and improving region recall represent a fundamental trade-off on this dataset, achieving a better balance will require a deeper structural understanding of abstract line art, highlighting an important direction for future work.
Matching Type AP Best F1 Precision Recall Top-1 Acc. Top-5 Acc.
Intra-img ($I_{a}$) 99.92 98.57 99.06 98.08
@@ B.2 Quantitative Results
Table C: Patch-level evaluation on PBC datasets (percentage values).
[…]
Intra-img ($I_{a}$) 83.05 60.20 20.30 0.32
@@
Cross-img 59.92
@@
Table D: Region-level evaluation on PBC datasets (percentage values; CR is shown in raw value). Region Precision and Region Recall are computed using predicted region pairs with purity >> 0.8.
Figure A: Example of region-level correspondence annotations in the PBC dataset. Matched regions are visualized using the same color across two images. Yellow boxes display the region ID and label ID of each region in the format Region ID (Label ID).
B.3 Qualitative Results
@@
Figure B: Visualization examples on the PBC dataset. In the region segmentation results, the same color within an image denotes a single region, and white indicates the background. In the region matching results, the same color across two images indicates a matched region pair, while gray denotes regions without a match.
@@
Following the qualitative analysis in Section 5.5 of the main paper, we evaluate our method on the PBC dataset from two aspects: (1) intra-image region segmentation and (2) cross-image region matching. Representative examples are shown in Figure B.
@@
For region segmentation, we compare our method with the Segment Anything Model (SAM) [kirillov2023segment]. As observed in Figure B, SAM often merges distinct semantic parts (e.g., face or clothing) with the background and fails to produce fine-grained boundaries due to the lack of color, texture, and shading cues in line art. In contrast, our method produces more precise and semantically coherent region groups.
For region matching, our approach accurately aligns major semantic regions, including cases with noticeable pose variation or shape deformation. The resulting region correspondences generally agree with human perception and demonstrate the applicability of our framework to clean, closed-contour line art.